OpenStudy (anonymous):

Quadratic Formula Can anyone explain the steps to this problem.. the only thing I need to ask is where does the "C" go if there is no C?

OpenStudy (anonymous):

2x^2-6=0

OpenStudy (anonymous):

if there is no C don't use the formula ok in your example there is no b

OpenStudy (anonymous):

much easier to write \[2x^2=6\\ x^2=3\\ x=\pm\sqrt3\] and be done no formula necessary

OpenStudy (anonymous):

Oh I see. Thanks so much! ;P

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Another question real quick.. If the problem were related someway to this one.. would I still have to leave it as it is? Just making sure.